OpenStudy (anonymous):

[\sqrt[5]{-1/3125}\]

OpenStudy (anonymous):

write it manually

OpenStudy (anonymous):

-1/5

OpenStudy (anonymous):

take the fifth root of both the top and the bottom, +5^5=3125 -1^5=-1

OpenStudy (anonymous):

\[\sqrt[5]{-1/3125}\]